# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

### v1

@@ 4.3.3. Adaptation to new corpora
The solution to the adaptation to new corpus challenge involves data augmentation, distributionally robust optimization, and domain-invariant projection.
@@
Data augmentation involves generating or modifying data in such a way that it bridges the gap between the source domain (the domain where the model was originally trained) and the target domain (the new domain where the model is to be applied). This can include techniques like synthesizing new data examples through transformations that maintain the integrity of the underlying patterns, translating examples from one domain to another, or creating semi-synthetic samples. The GPL (Wang et al., 2022b) uses an unsupervised domain adaptation method Generative Pseudo Labeling, which combines a query generator with pseudo labeling from a cross-encoder. The HyperR (Cai et al., 2022) performs a hyper-prompted training mechanism to enable uniform retrieval across tasks of different domains. There are other works that conduct unsupervised pre-training by using large-scale positive and negative pairs with different data augmentation methods, such as, query generation (Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023), synthetic pre-training (Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022), or synthetic relevance labels (Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b). Overall, data augmentation can enrich the training set to include more domain-relevant variations, thereby enhancing the model’s ability to generalize across domains.
@@
Domain modeling seeks to model the data from both the source and target domains into a common feature space where the differences between the domains are minimized. The idea is to learn a representation of the data that retains the essential information for the task at hand while discarding domain-specific features that might lead to bias or overfitting. By doing so, the model learns to focus on the underlying task without being distracted by differences between the domains. COCO-DR (Yu et al., 2022) use implicit distributionally robust optimization to reweight samples from different source query clusters for improving model robustness over rare queries during fine-tuning. Together with contrastive learning, this approach significantly improves the generalization of DRM over different corpora. There have been many successive efforts to optimize for this problem, including MoDIR (Xin et al., 2022), and ToTER (Kang et al., 2024). Xu et al. (2023) solved the domain OOD challenge by modeling a single passage as multiple units with two objectives. One is the semantic balance between units and the other is the extractability of essential units. Distributionally robust optimization helps in reducing the sensitivity of the model to changes in data distribution, thereby improving its adaptability.
Another way to deal with the new domain data is domain-invariant projection. Zhan et al. (2022) has been the first to use a relevance estimation module for modeling domain-invariant matching patterns and several domain adaption modules for modeling domain-specific features of multiple target corpora. Xian et al. (2023) propose a list-level alignment method, which aligns the distributions of the lists and preserves their list structure. They also demonstrate the superiority of the method on theoretical grounds. The domain-invariant feature space enables the model to perform well on the target domain using knowledge acquired from the source domain, thereby facilitating effective domain adaptation.
@@ 4.3.4. Updates to a corpus
In this scenario, IR models need to be constantly compatible with newly added documents, however, this can lead to catastrophic forgetting problems with old documents. Therefore, continuous learning (De Lange et al., 2021) has become the dominant approach, which aims at adapting the model to the newly added data without losing the ability to understand the old data by quickly adapting to the new unlabeled (little labeled) data.
@@
In GR, a sequence-to-sequence model is adopted to unify both the indexing and retrieval stages. All document information is encoded into the model parameters. The tight binding of the index to the retrieval module makes updating the index costly. To tackle this challenge, DSI++ (Mehta et al., 2023a) adapts a continual learning method for DSI (Tay et al., 2022) to incrementally index new documents while maintaining the ability to answer user queries related to both previously and newly indexed documents. After that, CorpusBrain++ (Guo et al., 2024) uses the continual learning method on another GR model called CorpusBrain (Chen et al., 2022b). Chen et al. (2023d) propose CLEVER to incrementally index new documents while supporting the ability to query both newly encountered documents and previously learned documents. CLEVER performs incremental product quantization (Jegou et al., 2010) to update a partial quantization codebook, and use a memory-augmented learning mechanism to form meaningful connections between old and new documents. Subsequent work on continuous learning has been devoted to the problem of guaranteeing to updates to a corpus for different GR models, for example, DynamicIR (Yoon et al., 2023) and IncDSI (Kishore et al., 2023).
@@
In DR, the model needs to learn the representation space of the entire corpus and encode each document into an embedding to serve as the index. Therefore, continual learning for DR should effectively adapt to the evolving distribution with the unlabeled new-coming documents, and avoid re-inferring all embeddings of old documents to efficiently update the index each time the model is updated. L2R (Cai et al., 2023) uses backward-compatible representations to deal with this problem. It first selects diverse support negatives for model training, and then uses a ranking alignment objective to ensure the backward-compatibility of representations.
4.4. OOD generalizability on unseen queries
In the dynamic landscape of information retrieval, where user queries constantly evolve, the ability of IR systems to generalize to unseen queries is paramount. Deep-learning based models, constrained by their training data, often falter when faced with novel query formulations. However, statistics indicate that 15% of daily Google searches are brand new. [^5] https://blog.google/products/search/our-latest-quality-improvements-search/ This challenge becomes particularly pronounced in the context of ever-changing user interests and the introduction of new terminologies. Previous work has analyzed the generalizability of existing IR models under query variants and different query types (Wu et al., 2022b; Penha et al., 2022; Zhuang et al., 2022; Liu et al., 2023b), respectively. In this work, we summarize existing work and how to improve the generalizability of IR models under unseen queries. The specific methodology categorization is shown in Figure 12.
OOD generalizability on unforeseen corpus Unseen query type (Cohen et al., 2018; Bigdeli et al., 2024; Ma et al., 2021a; Lupart et al., 2023; Sciavolino et al., 2021) Query variation Hybird methods (Tasawong et al., 2023; Campos et al., 2023; Sidiropoulos and Kanoulas, 2024; Pan et al., 2023) Contrastive learning methods (Zhuang and Zuccon, 2021; Liu et al., 2023a; Ma et al., 2021a; Sidiropoulos and Kanoulas, 2024) Self-teaching methods (Chen et al., 2022a; Zhuang and Zuccon, 2022; Zhuang et al., 2023)
@@ 4.4.1. Definition
Generalizability to unseen queries indicates the capacity of the IR model to sustain its retrieval performance when confronted with new and unfamiliar query formulations. In the field of IR, enhancing a model’s OOD generalizability with respect to unseen queries mainly involves improving the retrieval accuracy of the IR model across a variety of novel queries. Without loss of generality, given a test set $\tilde{\mathcal{D}}_{\mathrm{test}}$ comprising new queries $\tilde{\mathcal{Q}}$, which introduce a new distribution $\tilde{\mathcal{G}}$, the objective of augmenting the OOD generalizability of a neural IR model $f$ for unseen queries under top-$K$ ranked results can be formalized as:
@@
Query variations refer to different expressions of the same information need. The way in which information is expressed may impact the effectiveness of IR models (Penha et al., 2022). Some query variants, which introduce additional information, such as query expansion, tend to enhance the retrieval performance (Belkin et al., 1995; Benham et al., 2018). Some introduce noise, such as typos, grammatical errors, and variations in word order, which often challenges the robustness of the IR model (Penha et al., 2022; Campos et al., 2023; Zhuang and Zuccon, 2021). In this paper, we mainly focus on the latter one.
@@
(20) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};G\left(\mathcal{D}_{\mathrm{test}}\right),K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}\in\mathcal{D}_{\mathrm{train}},G\left(\mathcal{Q}\right)_{\sim\tilde{\mathcal{G}}}\in G\left(\mathcal{D}_{\mathrm{test}}\right).$
@@
(21) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{\tau_{i}}};\mathcal{D}_{\mathrm{test}}^{\tau_{j}},K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}^{\tau_{i}}\in\mathcal{D}_{\mathrm{train}}^{\tau_{i}},\mathcal{Q}_{\sim\mathcal{G}}^{\tau_{j}}\in\mathcal{D}_{\mathrm{test}}^{\tau_{j}}.$
@@
In the following, we first introduce evaluation metrics that are widely used for OOD generalizability on unseen queries. Then, we detail existing solutions for query variation and unseen query type, respectively.
@@ 4.4.3. Query variation
The solution to the query variation challenge involves (i) self-teaching method, (ii) contrastive learning method, and (iii) hybrid method.
@@
The self-teaching method focuses on distilling the matching capabilities of the IR model on the original clean query to the case of query variants. These methods often align the model output for distillation. Chen et al. (2022a) argue that the drift between query variations and original queries in model representation space affects the subsequent effectiveness of IR models. Based on this, they proposed RoDR, calibrating the in-batch local ranking of query variation to that of the original query for the representation space alignment. Zhuang and Zuccon (2022) also noticed this issue and employed CharacterBERT (El Boukkouri et al., 2020) as the backbone encoder to perform a character-level self-teaching method. This method distilled knowledge from queries without typos into the queries with typos in a character embedding space. ToRoDer (Zhuang et al., 2023) uses a pre-training method that uses bottlenecked information to recover the query variation.
@@
Contrastive learning methods typically make the model robust to query variants by enhancing the supervision of query variants against the original query as well as positive and negative samples. Zhuang and Zuccon (2021) propose a simple typos-aware training method for BERT-based DRMs and NRMs. During training, this method randomly selected query variants and migrated the supervised signals of the original queries directly for training. By comparing the similarity between a query and its variations and other distinct queries with contrastive learning, Sidiropoulos and Kanoulas (2022) improve the robustness of IR models when encountering typos. After that, MIRS (Liu et al., 2023a) uses a robust contrastive method via injection of [MASK] tokens into query variations and encouraging the representation similarity between the original query and the variation. There is also work on using contrastive learning to enhance the ability of the NRM to handle query variants (Ma et al., 2021a), as well as casting multi-positive contrastive learning (Sidiropoulos and Kanoulas, 2024).
@@
Some work enabled IR models to perform better than previously when dealing with query variants through a clever combination of self-teaching and contrastive learning. Tasawong et al. (2023) propose a typo-robust representation learning method that combines contrastive learning with dual self-teaching achieving competitive performance. CAPOT (Campos et al., 2023) introduces a notion of an anchoring loss between the unaltered model and the aligned model and designs a contrastive alignment post-training method to learn a robust model. Sidiropoulos and Kanoulas (2024) argue that previous work did not make full use of positive samples and employed contrastive learning with self-teaching that supports multiple positives. There is also work that uses LLMs to enhance the ability to deal with typos (Pan et al., 2023).
@@ 4.4.4. Unseen query type
Usually, part of the unseen query type is caused by unseen documents. Solutions to such problems are presented in Section 14. In this subsection, we present additional solutions for unseen query types.
Wu et al. (2022b) comprehensively analyze the robustness of existing ranking models in the face of unseen query types with 5 types of queries. They found that most NRMs do not generalize well to unseen query types. Even after training on multiple types of queries, NRMs still perform poorly when faced with a new kind of query. Among all ranking models, traditional probabilistic ranking models, such as BM25 (Robertson and Walker, 1994), have the strongest generalizability to OOD query types, while NRMs are the worst. Liu et al. (2023b) analyze the robustness of the generative retrieval model and the dense retrieval model under different query types and found that both models are sensitive to query types.
To improve the generalizability of unseen query types for IR models, Cohen et al. (2018) explore the use of adversarial learning as a regularization technique across different domains within the ranking task framework. By employing an adversarial discriminator and training the NRM across a limited number of domains, the discriminator acts to give negative feedback, thereby preventing the model from adopting domain-specific representations. Bigdeli et al. (2024) propose to integrate two knids of triplet loss functions into neural rankers such that they ensure that each query is moved along the embedding space, through the transformation of its embedding representation, in order to be placed close to its relevant document. In this way, they provided the opportunity to jointly rank documents and difficult queries. There has been some work focusing on the challenges of unseen query types in different scenarios (Ma et al., 2021a; Lupart et al., 2023; Sciavolino et al., 2021).
@@ 5. Open Issues and Future Directions
The robust neural IR presents several remaining challenges and opportunities for future research.
@@ 5.1. Remaining challenges and issues
In this subsection, we introduce the existing challenges and issues in robust neural IR that need to draw attention to adversarial robustness and OOD robustness, respectively.
@@ 5.1.1. Challenges on adversarial robustness in IR
Many other key issues in adversarial robustness in IR have not received much attention.
@@
Existing work focuses on attacking the first-stage retrieval (Liu et al., 2023c; Zhong et al., 2023) and re-rank stage (Liu et al., 2022a; Wu et al., 2023a) separately. Penetration attacks focus on exploiting vulnerabilities within the retrieval re-rank pipeline, a critical component in IR systems that ranks results based on relevance to the query. These attacks aim to manipulate rankings by identifying and exploiting weaknesses in the pipeline’s design or its underlying algorithms. This manipulation can result in irrelevant or malicious content being ranked higher than genuine content, compromising the integrity and reliability of the IR system.
@@
Universal attacks represent a form of adversarial threat that is particularly challenging due to its generalizability across different models and instances (Wallace et al., 2019). Unlike targeted attacks that aim at specific vulnerabilities within a system, universal attacks exploit common weaknesses that are present across a wide range of systems. This makes them incredibly difficult to defend against, as they require solutions that are not just effective for a single model or instance but across the entire spectrum of possible configurations. The development of robust defenses against universal attacks is therefore a significant challenge that demands innovative approaches and a deep understanding of the underlying vulnerabilities.
@@
Most existing work on adversarial attacks is based on static assumptions about search engines (Wu et al., 2023a; Liu et al., 2022a; Chen et al., 2023c). However, search engines operate within a dynamic landscape, which may include changes such as the expansion or reduction of the corpus, and the demotion of documents suspected of spamming in their rankings. In search engines, the search engine results page (SERP) for a query is constantly changing. Existing research indicates that, in the dynamic environment of search engines, current attack methods struggle to maintain a consistent ranking advantage (Liu et al., 2023d). Therefore, designing attack methods that fully consider the dynamic nature of search engines is both practically significant and challenging. At the same time, it’s also worthwhile to explore the development of defense methods that evolve in tandem with updates to the search engine.
Gaming in search engine
@@
In existing IR counter defenses, it is often the case that empirical defenses where the attack method is known, can yield good results (Liu et al., 2024b). Whereas in real scenarios, the attack methods are multiple and potentially unknown. Defending against unseen attacks is a paramount challenge in enhancing adversarial robustness in IR systems. These attacks are particularly daunting because they exploit new or unknown vulnerabilities, making traditional defense mechanisms, which are often designed to combat known threats, ineffective. The key to overcoming this challenge lies in the development of adaptive, intelligent systems capable of anticipating potential threats and dynamically adjusting their defense mechanisms in real-time. Achieving this level of adaptability and foresight requires a profound shift in the current paradigms of security in IR, embracing more proactive and predictive approaches.
Defense in practical search engines
Implementing effective defense mechanisms in practical search engines is a balancing act between effectiveness, efficiency, and cost. Effective defense strategies are those that can accurately detect and neutralize threats without significantly impacting the user experience or the relevance of search results. However, the computational resources required for these strategies often come with high costs and can affect the efficiency of the search engine, leading to slower response times and decreased user satisfaction. To address these challenges, search engines are increasingly turning to machine learning and artificial intelligence technologies that can provide scalable and cost-effective solutions. These technologies enable the development of adaptive defense mechanisms that can learn from attack patterns and evolve over time, offering a dynamic approach to security that maintains the delicate balance between protecting the search engine and preserving its performance.

### v2

@@ 4.3.3. Adaptation to new corpora
Proposed solutions to the adaptation to a new corpus involve data augmentation, distributionally robust optimization, and domain-invariant projection.
@@
Data augmentation involves generating or modifying data in such a way that it bridges the gap between the source domain (the domain where the model was originally trained) and the target domain (the new domain where the model is to be applied). This can include techniques like synthesizing new data examples through transformations that maintain the integrity of the underlying patterns, translating examples from one domain to another, or creating semi-synthetic samples. The GPL (Wang et al., 2022b) uses an unsupervised domain adaptation method generative pseudo labeling, which combines a query generator with pseudo labeling from a cross-encoder. HyperR (Cai et al., 2022) performs a hyper-prompted training mechanism to enable uniform retrieval across tasks of different domains.
There is other work that conducts unsupervised pre-training by using large-scale positive and negative pairs with different data augmentation methods such as, query generation (Liang et al., 2020; Ma et al., 2021b; Chandradevan et al., 2024; Thakur et al., 2021; Ma et al., 2021b; Sachan et al., 2023), synthetic pre-training (Reddy et al., 2021; Yu et al., 2022; Izacard et al., 2021; Neelakantan et al., 2022), or synthetic relevance labels (Li and Gaussier, 2024; Ge et al., 2023; Ram et al., 2022; Dai et al., 2022; Liu et al., 2022b). Overall, data augmentation can enrich the training set to include more domain-relevant variations, thereby enhancing the model’s ability to generalize across domains.
@@
Domain modeling seeks to model the data from both the source and target domains into a common feature space where the differences between the domains are minimized. The idea is to learn a representation of the data that retains the essential information for the task at hand while discarding domain-specific features that might lead to bias or overfitting. By doing so, the model learns to focus on the underlying task without being distracted by differences between the domains. COCO-DR (Yu et al., 2022) use implicit distributionally robust optimization to reweight samples from different source query clusters for improving model robustness over rare queries during fine-tuning. Together with contrastive learning, this approach significantly improves the generalization of DRM over different corpora. There have been many successive efforts to optimize for this problem, including MoDIR (Xin et al., 2022), and ToTER (Kang et al., 2024). Xu et al. (2023) address the domain OOD challenge by modeling a single passage as multiple units with two objectives. One is the semantic balance between units and the other is the extractability of essential units. Distributionally robust optimization helps in reducing the sensitivity of the model to changes in data distribution, thereby improving its adaptability.
Another way to deal with the new domain data is domain-invariant projection. Zhan et al. (2022) have been the first to use a relevance estimation module for modeling domain-invariant matching patterns and several domain adaption modules for modeling domain-specific features of multiple target corpora. Xian et al. (2023) propose a list-level alignment method, which aligns the distributions of the lists and preserves their list structure. They also demonstrate the superiority of their method on theoretical grounds. The domain-invariant feature space enables the model to perform well on the target domain using knowledge acquired from the source domain, thereby facilitating effective domain adaptation.
@@ 4.3.4. Updates to a corpus
In this scenario, IR models need to be compatible with newly added documents, however, this can lead to catastrophic forgetting problems with old documents. Therefore, continuous learning (De Lange et al., 2021) has become a dominant approach, which aims at adapting the model to the newly added data without losing the ability to understand the old data by quickly adapting to the new unlabeled (little labeled) data.
@@
In generative retrieval, a sequence-to-sequence model is adopted to unify both the indexing and retrieval stages. All document information is encoded into the model parameters. The tight binding of the index to the retrieval module makes updating the index costly. To tackle this challenge, DSI++ (Mehta et al., 2023a) adapts a continual learning method for DSI (Tay et al., 2022) to incrementally index new documents while maintaining the ability to answer user queries related to both previously and newly indexed documents. After that, CorpusBrain++ (Guo et al., 2024) uses a continual learning method on another generative retrieval model called CorpusBrain (Chen et al., 2022b). Chen et al. (2023d) propose CLEVER to incrementally index new documents while supporting the ability to query both newly encountered documents and previously learned documents. CLEVER performs incremental product quantization (Jegou et al., 2010) to update a partial quantization codebook, and use a memory-augmented learning mechanism to form meaningful connections between old and new documents. Subsequent work on continuous learning has been devoted to the problem of guaranteeing to updates to a corpus for different generative retrieval models, for example, DynamicIR (Yoon et al., 2023) and IncDSI (Kishore et al., 2023).
@@
In dense retrieval, the model needs to learn the representation space of the entire corpus and encode each document into an embedding to serve as the index. Therefore, continual learning for dense retrieval should effectively adapt to the evolving distribution with the unlabeled new-coming documents, and avoid re-inferring all embeddings of old documents to efficiently update the index each time the model is updated. L2R (Cai et al., 2023) uses backward-compatible representations to deal with this problem. It first selects diverse support negatives for model training, and then uses a ranking alignment objective to ensure the backward-compatibility of representations.
4.4. OOD Generalizability to Unseen Queries
Deep-learning based models, constrained by their training data, often falter when faced with novel query formulations. Previous work has analyzed the generalizability of IR models under query variants and different query types (Wu et al., 2022b; Penha et al., 2022; Zhuang et al., 2022; Liu et al., 2023b), respectively. Next, we summarize prior work and how to improve the generalizability of IR models under unseen queries. The specific methodology categorization is shown in Figure 11.
OOD generalizability on unforeseen corpus Unseen query type (Cohen et al., 2018; Bigdeli et al., 2024; Ma et al., 2021a; Lupart et al., 2023; Sciavolino et al., 2021) Query variation Hybrid methods (Tasawong et al., 2023; Campos et al., 2023; Sidiropoulos and Kanoulas, 2024; Pan et al., 2023) Contrastive learning methods (Zhuang and Zuccon, 2021; Liu et al., 2023a; Ma et al., 2021a) Self-teaching methods (Chen et al., 2022a; Zhuang and Zuccon, 2022; Zhuang et al., 2023)
@@ 4.4.1. Definition
Generalizability to unseen queries indicates the capacity of an IR model to sustain its retrieval performance when confronted with new and unfamiliar query formulations. Enhancing a model’s OOD generalizability with respect to unseen queries mainly involves improving the retrieval accuracy of the IR model across a variety of novel queries. Without loss of generality, given a test set $\tilde{\mathcal{D}}_{\mathrm{test}}$ comprising new queries $\tilde{\mathcal{Q}}$, which introduce a new distribution $\tilde{\mathcal{G}}$, the objective of augmenting the OOD generalizability of a neural IR model $f$ for unseen queries under top-$K$ ranked results can be formalized as:
@@
Query variations refer to different expressions of the same information need. The way in which information is expressed may impact the effectiveness of IR models (Penha et al., 2022). Some query variants, which introduce additional information, e.g., through query expansion, tend to enhance the retrieval performance (Belkin et al., 1995; Benham et al., 2018). Some introduce noise, such as typos, grammatical errors, and variations in word order, which often challenges the robustness of the IR model (Penha et al., 2022; Campos et al., 2023; Zhuang and Zuccon, 2021). In this paper, we mainly focus on the latter one.
@@
(20) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}};G\left(\mathcal{D}_{\mathrm{test}}\right),K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}\in\mathcal{D}_{\mathrm{train}},G\left(\mathcal{Q}\right)_{\sim\tilde{\mathcal{G}}}\in G\left(\mathcal{D}_{\mathrm{test}}\right),$
@@
(21) $\max\mathcal{R}_{M}\left(f_{\mathcal{D}_{\mathrm{train}}^{\tau_{i}}};\mathcal{D}_{\mathrm{test}}^{\tau_{j}},K\right)\text{ such that }\mathcal{Q}_{\sim\mathcal{G}}^{\tau_{i}}\in\mathcal{D}_{\mathrm{train}}^{\tau_{i}},\mathcal{Q}_{\sim\mathcal{G}}^{\tau_{j}}\in\mathcal{D}_{\mathrm{test}}^{\tau_{j}},$
@@
Next, we first introduce evaluation metrics that are widely used for OOD generalizability on unseen queries. Then, we detail existing solutions for query variation and unseen query type, respectively.
@@ 4.4.3. Query variation
Solutions to the query variation challenge include (i) a self-teaching method, (ii) a contrastive learning method, and (iii) a hybrid method.
@@
The self-teaching approach to query variations focuses on distilling the matching capabilities of the IR model on the original clean query to the case of query variants. These methods often align the model output for distillation. Chen et al. (2022a) argue that the drift between query variations and original queries in model representation space affects the subsequent effectiveness of IR models. Based on this, they propose RoDR, which calibrates the in-batch local ranking of query variants to that of the original query for the representation space alignment. Zhuang and Zuccon (2022) also notice this issue and employ CharacterBERT (El Boukkouri et al., 2020) as the backbone encoder to perform a character-level self-teaching method. This method distills knowledge from queries without typos into the queries with typos in a character embedding space. ToRoDer (Zhuang et al., 2023) uses a pre-training method that uses bottlenecked information to recover the query variation.
@@
Contrastive learning-based approaches to query variation typically make the model robust to query variants by enhancing the supervision of query variants against the original query as well as positive and negative samples. Zhuang and Zuccon (2021) propose a simple typos-aware training method for BERT-based DRMs and NRMs. During training, this method randomly selects query variants and migrated the supervised signals of the original queries directly for training. By comparing the similarity between a query and its variations and other distinct queries with contrastive learning, Sidiropoulos and Kanoulas (2022) improve the robustness of IR models when encountering typos. MIRS (Liu et al., 2023a) uses a robust contrastive method via injection of [MASK] tokens into query variations and encouraging the representation similarity between the original query and the variation.
@@
Some work enables IR models to perform better than previously when dealing with query variants through a combination of self-teaching and contrastive learning. Tasawong et al. (2023) propose a typo-robust representation learning method that combines contrastive learning with dual self-teaching achieving competitive performance. CAPOT (Campos et al., 2023) introduces a notion of an anchoring loss between the unaltered model and the aligned model and designs a contrastive alignment post-training method to learn a robust model. Sidiropoulos and Kanoulas (2024) argue that previous work does not make full use of positive samples and employ contrastive learning with self-teaching that supports multiple positives. There is also work that uses LLMs to enhance the ability to deal with typos (Pan et al., 2023).
@@ 4.4.4. Unseen query type
Usually, part of the unseen query type is caused by unseen documents. Solutions to such problems are presented in listed in Section 4.3. In this subsection, we present additional solutions for unseen query types.
Wu et al. (2022b) analyze the robustness of existing ranking models in the face of unseen query types with five types of queries. They find that most NRMs do not generalize well to unseen query types. Even after training on multiple types of queries, NRMs still perform poorly when faced with a new kind of query. Among all ranking models, traditional probabilistic ranking models, such as BM25 (Robertson and Walker, 1994), have the strongest generalizability to OOD query types, while NRMs are the worst. Liu et al. (2023b) analyze the robustness of the generative retrieval model and the dense retrieval model under different query types and find that both models are sensitive to query types.
To improve the generalizability to unseen query types for IR models, Cohen et al. (2018) explore the use of adversarial learning as a regularization technique across different domains within the ranking task framework. By employing an adversarial discriminator and training a NRM across a limited number of domains, the discriminator acts to give negative feedback, thereby preventing the model from adopting domain-specific representations. Bigdeli et al. (2024) propose to integrate two kinds of triplet loss functions into neural rankers such that they ensure that each query is moved along the embedding space, through the transformation of its embedding representation, in order to be placed close to its relevant document. In this way, they provide the opportunity to jointly rank documents and difficult queries. There has been some work focusing on the challenges of unseen query types in different scenarios (Ma et al., 2021a; Lupart et al., 2023; Sciavolino et al., 2021).
@@ 5. Open Issues and Future Directions
In addition to the significant progres documented in Section 3 and 4 above, robust neural IR presents several remaining challenges and opportunities for future research.
@@ 5.1. Remaining Challenges and Issues
In this subsection, we identify challenges and issues in robust neural IR with a special focus on adversarial robustness and OOD robustness, respectively.
@@ 5.1.1. Challenges on adversarial robustness in IR
Many key issues in adversarial robustness in IR have not received much attention.
@@
As explained in Section 3 and 4, there is a considerable body of work that focuses on attacking the first-stage retrieval (Liu et al., 2023c; Zhong et al., 2023) and re-rank stage (Liu et al., 2022a; Wu et al., 2023a) separately. Penetration attacks focus on exploiting vulnerabilities within the retrieval re-rank pipeline, a critical component in IR systems that ranks results based on relevance to the query. These attacks aim to manipulate rankings by identifying and exploiting weaknesses in the pipeline’s design or its underlying algorithms. This manipulation can result in irrelevant or malicious content being ranked higher than genuine content, compromising the integrity and reliability of the IR system.
@@
Universal attacks represent a form of adversarial threat that is particularly challenging due to its generalizability across different models and instances (Wallace et al., 2019). Unlike targeted attacks that aim at specific vulnerabilities within a system, universal attacks exploit common weaknesses that are present across a wide range of systems. This makes it difficult to defend against them, as they require solutions that are not just effective for a single model or instance but across the entire spectrum of possible configurations. The development of robust defenses against universal attacks is therefore a significant challenge that demands innovative approaches and a deep understanding of the underlying vulnerabilities.
@@
Most prior work on adversarial attacks is based on static assumptions about search engines (Wu et al., 2023a; Liu et al., 2022a; Chen et al., 2023c). However, search engines operate within a dynamic landscape, which may include changes such as the expansion or reduction of the corpus, and the demotion of documents suspected of spamming in their rankings. In search engines, the search engine results page (SERP) for a query is constantly changing. Research indicates that, in the dynamic environment of search engines, current attack methods struggle to maintain a consistent ranking advantage (Liu et al., 2023d). Therefore, designing attack methods that fully consider the dynamic nature of search engines is both practically significant and challenging. At the same time, it is worthwhile to explore the development of defense methods that evolve in tandem with updates to the search engine.
Gaming in search engines
@@
In IR counter defenses, it is often the case that empirical defenses where the attack method is known, can yield good results (Liu et al., 2024b). Whereas in real scenarios, the attack methods are multiple and potentially unknown. Defending against unseen attacks is a paramount challenge in enhancing adversarial robustness in IR systems. These attacks are particularly daunting because they exploit new or unknown vulnerabilities, making traditional defense mechanisms, which are often designed to combat known threats, ineffective. The key to overcoming this challenge lies in the development of adaptive, intelligent systems capable of anticipating potential threats and dynamically adjusting their defense mechanisms in real-time. Achieving this level of adaptability and foresight requires a profound shift in the current paradigms of security in IR, embracing more proactive and predictive approaches.
Defense in practice
Implementing effective defense mechanisms in practice is a balancing act between effectiveness, efficiency, and cost. Effective defense strategies are those that can accurately detect and neutralize threats without significantly impacting the user experience or the relevance of search results. However, the computational resources required for these strategies often come with high costs and can affect the efficiency of the search engine, leading to slower response times and decreased user satisfaction. To address these challenges, search engines are increasingly turning to machine learning and artificial intelligence technologies that can provide scalable and cost-effective solutions (Leonhardt et al., 2024; Ma et al., 2022a). These technologies enable the development of adaptive defense mechanisms that can learn from attack patterns and evolve over time, offering a dynamic approach to security that maintains the delicate balance between protecting the search engine and preserving its performance.